OpenStudy (anonymous):

Estimate 2 and six-sevenths times 5 and one-fourth by rounding each factor to the nearest whole number. Explain your choices for rounding

OpenStudy (anonymous):

2 6/7 is nearly3 5 1/4 is nearly 5 3 x 5 = 15 exact answer is 15, too (20/7)(21/4) = 420/28 = 15 neat